# Debate, Train, Evolve: Self-Evolution of Language Model Reasoning

### v1

@@ Abstract
Large language models (LLMs) have improved significantly in their reasoning through extensive training on massive datasets. However, relying solely on additional data for improvement is becoming increasingly impractical, highlighting the need for models to autonomously enhance their reasoning without external supervision. In this paper, we propose Debate, Train, Evolve (DTE), a novel ground truth-free training framework that uses multi-agent debate traces to evolve a single language model. We also introduce a new prompting strategy Reflect-Critique-Refine, to improve debate quality by explicitly instructing agents to critique and refine their reasoning. Extensive evaluations on five reasoning benchmarks with six open-weight models show that our DTE framework achieve substantial improvements, with an average accuracy gain of 8.92% on the challenging GSM-PLUS dataset. Furthermore, we observe strong cross-domain generalization, with an average accuracy gain of 5.8% on all other benchmarks, suggesting that our method captures general reasoning capabilities.
@@ 1 Introduction
Recent approaches attempt to overcome the data bottleneck by enabling models to generate and learn from synthetic data, which is generated by automatically expanding a small set of seed tasks into large synthetic instruction datasets Wang et al. (2022); Zeng et al. (2024). Other methods Madaan et al. (2023); Jiang et al. (2023); Gou et al. (2023); Peng et al. (2023); Zelikman et al. (2024); Costello et al. (2025) refine model-generated outputs through iterative self-feedback or preference optimization. Despite their effectiveness, these self-evolution strategies predominantly rely on judgments from a single model or a teacher-student configuration, often leading to confirmation bias and insufficient reasoning diversity.
@@
First, we conduct a large-scale empirical analysis of MAD using open-source models, where we identify limitations of the original MAD prompting approach, particularly in smaller models Du et al. (2023). To address this, we propose a Reflect-Critique-Refine (RCR) prompting strategy, which explicitly forces agents to identify, critique, and correct reasoning errors in both their own and peers’ answers. Second, using this prompting strategy, we build our DTE framework (Figure 1). Finally, we find that models with parameters suffer accuracy loss Srivastava et al. (2025) after second evolution round; our controlled study shows that the problem correlates with large temperature-induced variance and high KL divergence from the base policy. Lowering the sampling temperature from 0.7 to 0.3 cuts the KL drift by and recovers up to 76% of the lost performance, preventing catastrophic forgetting in smaller models without extra supervision.
Our experiments show significant gains in reasoning performance across multiple datasets. Specifically, our evolved models show an average accuracy improvement of 8.92% on the challenging GSM-PLUS dataset compared to their original versions. Moreover, our framework achieves notable cross-domain generalization, enhancing model performance across datasets not seen during training. These results confirm that our Debate, Train, Evolve method successfully distills multi-agent debate’s insights into efficient single-model inference, bridging the gap between computational efficiency and advanced reasoning capabilities.
@@ 2 Related Work
Du et al. (2023) first showed that letting several large models debate improves accuracy on maths, strategy, and factual QA without any new parameters. Later, Liang et al. (2023) highlighted the risk of degeneration-of-thought: a single agent quickly converges on one path, whereas a two-debater plus judge setup maintains diversity and outperforms GPT-4 on tricky arithmetic. RECONCILE (Chen et al., 2023) mixes agents from different model families, reaches consensus through confidence-weighted votes, and adds up to eleven points on seven reasoning benchmarks. Smit et al. (2023) shows that MAD beats sampling ensembles only after careful tuning. Finally, works like PREDICT (Park et al., 2024) apply multi-agent debate to tasks beyond QA, such as hate-speech classification, where agents reason under different guidelines. Recent advances further incorporate explicit reinforcement learning into the debate process. For example, the ACC-Collab framework Estornell et al. (2024) utilized an actor-critic approach to explicitly optimize agent collaboration, yielding superior performance on reasoning tasks.
@@
SELF-INSTRUCT (Wang et al., 2022) prompts GPT-3 to write 52000 novel instructions plus answers and then fine-tunes on its own output, reducing the gap to InstructGPT by thirty-three points on Super-Natural-Instructions without extra human labels. STaR (Zelikman et al., 2024) augments a few chain-of-thought exemplars by letting the model explain wrong answers in reverse, doubling CommonsenseQA accuracy for a 350M model. SELF-REFINE (Madaan et al., 2023) and the broader SELF framework (Lu et al., 2023) turn one model into writer, critic and re-writer, looping feedback at inference or during fine-tuning to improve on GSM8K by around seven points. Instruction-tuning variants refine the idea: Self-Refine Instruction-Tuning (Ranaldi and Freitas, 2024) pairs Llama-2 and Mistral students with large teacher rationales and then lets each student prefer its own better reasoning, closing the size gap on commonsense and math tasks. More recently, Think, Prune, Train, Improve (Costello et al., 2025) shows that careful filtering of self-generated traces can raise Gemma-2B to 58% on GSM8K and push Llama-3-70B beyond GPT-4o. These studies confirm that single-agent loops, with or without ground truth, can expand a model’s ability.
@@
3 Method
Preliminary Study: Multi-Agent Debate (MAD)
First, we conduct a large-scale empirical study on MAD. Our initial objective was to understand how MAD affects performance across different reasoning tasks (eg., math, science), model families, and across scales, and to identify its limitations. Our ultimate goal was to generate high-quality reasoning traces using MAD. Following MAD Du et al. (2023) prompting strategy, however, our initial experiments show that the gains from MAD were minimal across reasoning benchmarks.
Our manual inspection uncovered two primary issues: (i) sycophancy: agents often discarded their correct answer and copied a other agent’s wrong but confident solution; (ii) verbosity bias: when two answers disagreed, agents preferred the longer rationale even if it was logically weak Saito et al. (2023). For example, on average, a three-agent Qwen-1.5B committee, sycophancy reached (28% of the debates have sycophancy rounds). These effects lead to a substantial fraction of [correct incorrect] transitions during debate.
3.1 Reflect–Critique–Refine Prompting
Inspired by previous works that had iterative self-feedback procedures Madaan et al. (2023); Gou et al. (2023); Peng et al. (2023), we propose Reflect-Critique-Refine (RCR) prompting, which mitigates this problem. Unlike the original MAD prompt that simply asks agents to revise their answer by looking at others, RCR prompting encourages agents to present confident reasoning. 1) Every agent must reflect by stating why its current answer could be wrong; 2) Each agent then critiques exactly two peer rationales, naming a specific flaw or a step it cannot verify; 3) Finally the agent may refine its answer—but only if it supplies at least one new reasoning step.
@@ 3.1 Reflect–Critique–Refine Prompting
Phrases like “identify any errors” reliably trigger negative tokens (“error”, “mistake”, “step X is wrong”) which LLMs have learned during supervised finetuning. By specifying valid next moves (defend/correct/adopt), we implicitly shape the log-probability mass toward useful trajectories, shrinking the space of rambling answers. The single-step explanation requirement forces agents to think before copying, while the fixed quota of peer critiques limits needless verbosity. This prompting helped reduce sycophancy and improved performance.
Debate protocol
Let be a set of agents and a query. Round 0: each produces , an answer and a rationale. If consensus (all answers identical) occurs, the process stops. Otherwise we enter a debate round following Algorithm 1: agents receive all previous rationales, execute Reflect–Critique–Refine, and output . The debate ends when either (i) consensus is reached or (ii) a maximum of rounds is exhausted, after which the final answer is decided by simple majority vote.
3.2 Debate, Train, Evolve Framework
So far we have seen that RCR prompting yields high-quality debate traces and measurable accuracy gains. We now move from learning from these debate traces to train a single language model. This results in Debate, Train, Evolve (DTE), a closed loop that distills the collective insight of multiple agents into a single, faster model.
Framework overview
Let be a frozen base policy with parameters . At iteration we run Algorithm 1 on a batch of queries , producing a set of debate traces where is the original query, the consensus answer and a compressed rationale extracted from the highest-scoring agent. Extraction step keeps any explanation steps that appear in at least two agents or that introduce a symbolic manipulation absent from the previous round.
We then fine-tune a student policy on with Group Relative Policy Optimization Shao et al. (2024) and then place back into the agent pool, replacing its earlier version. The loop repeats until the mean validation reward stagnates for two iterations or a maximum of five cycles is reached. Algorithm 2 gives the exact procedure.
Reward shaping and GRPO
For a query and a student output we define the shaped reward
where is the indicator function, checks the <reasoning> <answer> XML template, and is the token length.
Weights are set to and .
GRPO maximises the clipped surrogate
where , , and . The first term is identical to PPO but uses the debate-derived reward; the second anchors the student to the base policy, a key factor for avoiding catastrophic forgetting Luo et al. (2025); Kotha et al. (2024). Gradients are estimated with the standard REINFORCE algorithm and a value network baseline.
| Model | GSM8K | GSM-Plus | ARC-Challenge | ||||||
| Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | |
| Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | |
| Qwen-2.5-1.5B | 62.77 | 72.33 | 73.09 (+10.32 ) | 42.00 | 53.33 | 55.92 (+13.92 ) | 69.21 | 68.52 | 68.36 (-0.85 ) |
| Qwen-2.5-3B | 84.08 | 85.14 | 86.05 (+1.97 ) | 61.75 | 68.00 | 69.50 (+7.75 ) | 83.53 | 84.64 | 83.95 (-0.42 ) |
| Qwen-2.5-7B | 90.67 | 91.21 | 88.32 (-2.35 ) | 68.62 | 74.17 | 74.71 (+6.09 ) | 87.22 | 91.64 | 90.89 (+3.67 ) |
| Qwen-2.5-14B | 92.80 | 93.33 | 93.74 (+0.94 ) | 71.79 | 77.25 | 78.88 (+7.09 ) | 90.27 | 93.77 | 93.13 (+2.86 ) |
| Llama-3.2-3B | 72.55 | 73.84 | 75.06 (+2.51 ) | 45.67 | 51.12 | 53.79 (+8.12 ) | 73.12 | 76.19 | 77.23 (+4.11 ) |
| Llama-3.1-8B | 81.73 | 82.18 | 86.81 (+5.08 ) | 55.62 | 60.79 | 66.17 (+10.55 ) | 77.65 | 85.07 | 86.53 (+8.88 ) |
@@ 4.1 Experimental Setup
We conduct of RCR prompting study on ten open-weight models—Qwen (0.5-32B), Llama-3/8B, Mistral-7B, Phi-mini—and two proprietary models, GPT-4o and GPT-4o-mini. We study our DTE framework with 6 models (Qwen 1.5B-14B, Llama-3B and Llama-8B). Baselines are: (i) the single original model; (ii) vanilla MAD with the original MAD prompt Du et al. (2023).
@@
During debate we sample each agent once per query at temperature (exploratory) or (deterministic); mixed-teams use one exploratory and two deterministic agents. For evolution we adopt LoRA fine-tuning (rank 128, dropout 0.05) on attention and MLP projections, freezing embeddings and layer norms. GRPO is optimized with AdamW (learning rate , weight decay 0.01, 50-step linear warm-up). Each evolution epoch processes 8k debate traces (2 M tokens) and runs on A100-80 GB GPUs for a 7B model; larger models scale near-linearly.
@@
Task performance is exact match for GSM-style datasets and accuracy for MC-QA. For RCR evaluation, we also track Sycophancy-Rate: the fraction of agents switching to an incorrect peer answer without adding new reasoning; [incorrect correct] instances during MAD.

### v2

@@ Abstract
Large language models (LLMs) have improved significantly in their reasoning through extensive training on massive datasets. However, relying solely on additional data for improvement is becoming increasingly impractical, highlighting the need for models to autonomously enhance their reasoning without external supervision. In this paper, we propose Debate, Train, Evolve (DTE), a novel ground truth-free training framework that uses multi-agent debate traces to evolve a single language model. We also introduce a new prompting strategy Reflect-Critique-Refine, to improve debate quality by explicitly instructing agents to critique and refine their reasoning. Extensive evaluations on seven reasoning benchmarks with six open-weight models show that our DTE framework achieve substantial improvements, with an average accuracy gain of 8.92% on the GSM-PLUS dataset. Furthermore, we observe strong cross-domain generalization, with an average accuracy gain of 5.8% on all other benchmarks, suggesting that our method captures general reasoning capabilities. Our framework code and trained models are publicly available at https://github.com/ctrl-gaurav/Debate-Train-Evolve. 111GitHub: github.com/ctrl-gaurav/Debate-Train-Evolve DTE Website: https://ctrl-gaurav.github.io/debate-train-evolve.github.io/
Debate, Train, Evolve: Self-Evolution of Language Model Reasoning
Gaurav Srivastava, Zhenyu Bi, Meng Lu, Xuan Wang††thanks: Corresponding Author Department of Computer Science, Virginia Tech, USA (gks, zhenyub, menglu, xuanw)@vt.edu Website: debate-train-evolve.github.io/
@@ 1 Introduction
Recent approaches attempt to overcome the data bottleneck by enabling models to generate and learn from synthetic data, which is generated by automatically expanding a small set of seed tasks into large synthetic instruction datasets Wang et al. (2023); Zeng et al. (2024). Other methods Madaan et al. (2023); Jiang et al. (2023); Gou et al. (2023); Zelikman et al. (2024); Costello et al. (2025) refine model-generated outputs through iterative self-feedback or preference optimization. Despite their effectiveness, these self-evolution strategies predominantly rely on judgments from a single model or a teacher-student configuration, often leading to confirmation bias and insufficient reasoning diversity.
@@
First, we conduct a large-scale empirical analysis of MAD using open-source models, where we identify limitations of the original MAD prompting approach, particularly in smaller models Du et al. (2023). To address this, we propose a Reflect-Critique-Refine (RCR) prompting strategy, which explicitly forces agents to identify, critique, and correct reasoning errors in both their own and peers’ answers. Second, using this prompting strategy, we build our DTE framework (Figure 1). Finally, we find that models with parameters suffer accuracy loss Srivastava et al. (2025a, c, b) after second evolution round; our controlled study shows that the problem correlates with large temperature-induced variance and high KL divergence from the base policy. Lowering the sampling temperature from 0.7 to 0.3 cuts the KL drift by and recovers up to 76% of the lost performance, preventing catastrophic forgetting in smaller models without extra supervision.
Our experiments show significant gains in reasoning performance across multiple datasets. Specifically, our evolved models show an average accuracy improvement of 8.92% on GSM-PLUS dataset compared to their original versions. Moreover, our framework achieves notable cross-domain generalization, enhancing model performance across datasets not seen during training. These results confirm that our method successfully distills multi-agent debate’s insights into efficient single-model inference, bridging the gap between computational efficiency and improved reasoning.
@@ 2 Related Work
Du et al. (2023) first showed that letting several large models debate improves accuracy on maths, strategy, and factual QA without any new parameters. Later, Liang et al. (2024) highlighted the risk of degeneration‑of‑thought: a single agent quickly converges on one path, whereas a two‑debater plus judge setup maintains diversity and outperforms GPT‑4 on tricky arithmetic. RECONCILE (Chen et al., 2024) mixes agents from different model families, reaches consensus through confidence‑weighted votes, and adds up to eleven points on seven reasoning benchmarks. Smit et al. (2023) shows that MAD beats sampling ensembles only after careful tuning. Finally, works like PREDICT (Park et al., 2024) apply multi-agent debate to tasks beyond QA, such as hate-speech classification, where agents reason under different guidelines. Recent advances further incorporate explicit reinforcement learning into the debate process. For example, the ACC-Collab framework Estornell et al. (2024) utilized an actor-critic approach to explicitly optimize agent collaboration, yielding superior performance on reasoning tasks.
@@
SELF‑INSTRUCT (Wang et al., 2023) prompts GPT‑3 to write 52000 novel instructions plus answers and then fine‑tunes on its own output, reducing the gap to InstructGPT by thirty‑three points on Super‑Natural‑Instructions without extra human labels. STaR (Zelikman et al., 2024) augments a few chain‑of‑thought exemplars by letting the model explain wrong answers in reverse, doubling CommonsenseQA accuracy for a 350M model. SELF‑REFINE (Madaan et al., 2023) and the broader SELF framework (Lu et al., 2023) turn one model into writer, critic and re‑writer, looping feedback at inference or during fine‑tuning to improve on GSM8K by around seven points. Instruction‑tuning variants refine the idea: Self‑Refine Instruction‑Tuning (Ranaldi and Freitas, 2024) pairs Llama‑2 and Mistral students with large teacher rationales and then lets each student prefer its own better reasoning, closing the size gap on commonsense and math tasks. More recently, Think, Prune, Train, Improve (Costello et al., 2025) shows that careful filtering of self‑generated traces can raise Gemma‑2B to 58% on GSM8K and push Llama‑3‑70B beyond GPT‑4o. These studies confirm that single‑agent loops, with or without ground truth, can expand a model’s ability.
@@
3 Debate, Train, Evolve Framework
In this section, we first analyze limitations of existing multi-agent debate approaches (§3.1), introduce our improved prompting strategy (§3.2), and then detail the mathematical framework for training models using debate-derived rewards (§3.3). Our DTE framework uses multi-agent debate to generate high-quality reasoning traces, then distills these traces into a single model through group-relative policy optimization Shao et al. (2024).
3.1 Preliminary Analysis of Multi-Agent Debate
Let denote a set of language model agents, and let represent an input query. In the standard multi-agent debate framework, each agent independently generates an initial response consisting of an answer and rationale . Agents then engage in rounds of debate, where in round , each agent observes peer responses and produces an updated response .
Our empirical analysis of this standard approach revealed two critical failure modes. First, we observed high rates of sycophancy, where agents abandon correct answers in favor of incorrect but confidently-stated peer solutions. Second, we identified a verbosity bias where agents preferentially adopt longer rationales regardless of logical validity Saito et al. (2023). These effects resulted in degraded debate quality (substantial fraction of [correct incorrect] transitions during debate), particularly for smaller models where sycophancy rates exceeded 28% on average.
3.2 Reflect-Critique-Refine Prompting Strategy
To address these limitations, we introduce the RCR prompting strategy. Unlike standard debate prompts that simply request answer revision Madaan et al. (2023); Gou et al. (2023); Peng et al. (2023), RCR structures agent responses through three explicit phases: 1) Reflect: Each agent must identify potential errors in its current reasoning by generating a self-critique . 2) Critique: The agent then evaluates exactly two peer rationales, producing critiques where and . 3) Refine: Finally, the agent updates its response to subject to the constraint that if , then must contain at least one novel reasoning step not present in .
@@ 3.2 Reflect-Critique-Refine Prompting Strategy
Phrases like “identify any errors” reliably trigger negative tokens (“error”, “mistake”, “step X is wrong”) which LLMs have learned during supervised finetuning. By specifying valid next moves (defend/correct/adopt), we implicitly shape the log‑probability mass toward useful trajectories, shrinking the space of rambling answers. The single‑step explanation requirement forces agents to think before copying and reduces sycophancy by requiring agents to justify answer changes with novel reasoning, while the fixed critique quota prevents unbounded verbosity. Algorithm 1 presents the complete debate protocol, where the debate terminates when either consensus is reached (all identical) or after rounds, with the final answer determined by majority vote.
3.3 Training via Group Relative Policy Optimization
We now formalize how debate traces are used to train a single language model. Let denote a language model policy parameterized by , which models the conditional distribution over token sequences: , where is the input state (query) and is the generated token sequence.
Debate Trace Extraction and Reward Design
Given a query , we run multi-agent debate using Algorithm 1 to obtain a consensus answer and a set of reasoning traces . From these traces, we extract a consolidated rationale by identifying reasoning steps that either (i) appear in multiple agents’ responses or (ii) introduce novel symbolic manipulations. This yields a training instance . For each generated response to query , we define a shaped reward function:
where indicates answer correctness (verified via exact string match after normalization), checks adherence to the XML template structure, denotes token length, and with .
Group Relative Advantage Estimation
For training, we use Group Relative Policy Optimization (GRPO), which eliminates the need for a separate value function by estimating advantages through group-wise comparisons. For each query in our training batch, we sample responses from the current policy . Each response receives a scalar reward .
Instead of learning a value function to estimate expected returns, GRPO computes advantages using the group statistics. The advantage for response at token position is:
where is the mean reward, is the standard deviation, and prevents division by zero.
This formulation provides several key benefits. First, responses with above-average rewards receive positive advantages, encouraging the model to increase their likelihood. Second, normalization by standard deviation ensures that advantages remain stable across different reward scales. Third, using group statistics rather than a learned baseline reduces memory requirements by eliminating the value network.
Policy Optimization Objective
Given the group-relative advantages, we optimize the policy using a clipped surrogate objective with KL regularization. The GRPO loss for a single query is:
where the clipped policy gradient loss is:
Here, is the importance ratio between the new and old policies, and is the clipping threshold. The clipping mechanism prevents destructively large policy updates: when exceeds or falls below , the gradient contribution is capped.
The KL divergence term regularizes the policy to prevent excessive deviation from a reference model (typically the initial supervised fine-tuned model):
with regularization strength . This KL penalty serves a different purpose than the clipping: while clipping prevents large single-step updates, the KL term anchors the policy to maintain linguistic coherence and prevent catastrophic forgetting.
Gradient Estimation and Optimization
Gradient of with respect to is estimated using the REINFORCE algorithm. For each token in response , the gradient contribution is:
where is the effective advantage after clipping and KL regularization. This expectation is approximated through Monte Carlo sampling using the generated responses. We optimize using AdamW with learning rate , weight decay , and a 50-step linear warmup. To enhance training efficiency, we use LoRA (Low-Rank Adaptation) with rank and dropout probability , applying adaptations to attention and MLP projection matrices while keeping embeddings and layer normalizations frozen.
3.4 Evolution through Iterative Training
The complete DTE framework operates as an iterative process, formalized in Algorithm 2. Starting with a base policy , we perform evolution rounds where each round consists of: 1) Debate Generation: Sample a batch of queries and generate debate traces using RCR-prompted multi-agent debate (Algorithm 1), producing dataset . 2) Policy Update: Fine-tune on using GRPO to obtain . 3) Agent Replacement: Replace the previous version in the debate ensemble with the evolved policy.
The process continues until validation performance plateaus or a maximum number of iterations is reached. For smaller models (B parameters), we implement temperature annealing from to across rounds to mitigate KL divergence growth and prevent catastrophic forgetting, as high-temperature sampling in later rounds can cause excessive policy drift.
This framework achieves autonomous reasoning improvement by combining the exploration benefits of multi-agent debate with the efficiency of single-model deployment, while GRPO’s group-relative formulation provides stable training without requiring auxiliary value networks.
| Model | GSM8K | GSM-Plus | MATH | ARC-Challenge | GPQA Main | ||||||||||
| Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | Original | 3 Agent | Evolved Single | |
| Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | Model | MAD | Model (DTE) | |
| Qwen-2.5-1.5B | 62.77 | 72.33 | 73.09 (+10.32 ) | 42.00 | 53.33 | 55.92 (+13.92 ) | 45.08 | 50.68 | 52.20 (+7.12 ) | 69.21 | 68.52 | 68.36 (-0.85 ) | 19.42 | 18.75 | 20.10 (+0.68 ) |
| Qwen-2.5-3B | 84.08 | 85.14 | 86.05 (+1.97 ) | 61.75 | 68.00 | 69.50 (+7.75 ) | 61.36 | 65.72 | 67.10 (+5.74 ) | 83.53 | 84.64 | 83.95 (-0.42 ) | 28.12 | 29.24 | 30.50 (+2.38 ) |
| Qwen-2.5-7B | 90.67 | 91.21 | 88.32 (-2.35 ) | 68.62 | 74.17 | 74.71 (+6.09 ) | 73.08 | 75.58 | 77.20 (+4.12 ) | 87.22 | 91.64 | 90.89 (+3.67 ) | 32.81 | 33.71 | 35.20 (+2.39 ) |
| Qwen-2.5-14B | 92.80 | 93.33 | 93.74 (+0.94 ) | 71.79 | 77.25 | 78.88 (+7.09 ) | 76.18 | 78.62 | 80.10 (+3.92 ) | 90.27 | 93.77 | 93.13 (+2.86 ) | 41.29 | 42.19 | 43.60 (+2.31 ) |
| Llama-3.2-3B | 72.55 | 73.84 | 75.06 (+2.51 ) | 45.67 | 51.12 | 53.79 (+8.12 ) | 39.76 | 41.90 | 43.80 (+4.04 ) | 73.12 | 76.19 | 77.23 (+4.11 ) | 26.12 | 29.24 | 30.80 (+4.68 ) |
| Llama-3.1-8B | 81.73 | 82.18 | 86.81 (+5.08 ) | 55.62 | 60.79 | 66.17 (+10.55 ) | 46.66 | 47.90 | 49.40 (+2.74 ) | 77.65 | 85.07 | 86.53 (+8.88 ) | 27.46 | 32.37 | 34.10 (+6.64 ) |
@@ 4.1 Experimental Setup
We conduct experiments on seven public reasoning benchmarks: 1) GSM8K Cobbe et al. (2021), 2) GSM-Plus Li et al. (2024) (adversarial math problems), 3) MATH Hendrycks et al. (2021) (competition-level mathematics), 4) ARC-Easy, 5) ARC-Challenge (Clark et al., 2018) (science reasoning), 6) GPQA Main Rein et al. (2024) (graduate-level STEM questions), and 7) CommonsenseQA (Talmor et al., 2019).
@@
We conduct of RCR prompting study on ten open-weight models, Qwen (0.5-32B), Llama-3/8B, Mistral-7B, Phi-mini, and two proprietary models, GPT-4o and GPT-4o-mini. We study our DTE framework with 6 models (Qwen 1.5B-14B, Llama-3B and Llama-8B). Baselines are: (i) the single original model; (ii) vanilla MAD with the original MAD prompt.
@@
During debate we sample each agent once per query at temperature (exploratory) or (deterministic); mixed-teams use one exploratory and two deterministic agents. For GRPO training, we adopt LoRA fine-tuning (rank 128, dropout 0.05) on attention and MLP projections, freezing embeddings and layer norms. GRPO is optimized with AdamW (learning rate , weight decay , and momentum coefficients , ). We set the GRPO-specific hyperparameters as: clipping threshold , KL coefficient , and group size responses per query. Each evolution epoch processes 8k debate traces (2 M tokens) and runs on A100-80 GB GPUs for a 7B model; larger models scale near-linearly.
@@
Task performance is exact match for GSM-style datasets and accuracy for MC-QA. For RCR evaluation, we also track Sycophancy-Rate and [incorrect correct] instances.